OpenStudy (firejay5):

State the coordinates of the focus and the directrix equation for each parabola. *Show work and explain how you got what you got*

OpenStudy (firejay5):

\[x = \frac{ 1 }{ 20 } (y - 1)^2 - 4\]

OpenStudy (anonymous):

vertex is pretty clear right?

OpenStudy (firejay5):

yea (-4, 1)

OpenStudy (firejay5):

and the AOS = y = 1 right

OpenStudy (anonymous):

yeah that looks good

OpenStudy (anonymous):

and you know which way it faces right?

OpenStudy (anonymous):

not sure what you mean by "horizonal" since the y is squared it opens to the right or left, and since \(\frac{1}{20}\) is positive if opens to the right 

OpenStudy (firejay5):

How do you find the focus and directrix?

OpenStudy (anonymous):

it is easier if we write it as \[20(x+4)=(y-1)^2\] then we see that it looks like \[4p(x-h)=(y-k)^2\] making \(4p=20\) and so \(p=5\) then we can find the focus and the directrix easily

OpenStudy (firejay5):

Well I did it this way: \[\frac{ 1 }{ 4(\frac{ 1 }{ 20 }) } = 5\]

OpenStudy (anonymous):

ok that works too

OpenStudy (firejay5):

that's the way we use it in class

OpenStudy (anonymous):

so you know \(p=5\) and that makes the focus 5 units to the right of \((-4,1)\) and the directrix 5 units to the left

OpenStudy (anonymous):

that is why you have to know which way it is oriented, to know whether to go up and down or left and right

OpenStudy (firejay5):

do we mess with the x or y coordinate

OpenStudy (anonymous):

since it is right and left, mess with the x's

OpenStudy (firejay5):

I know you would add 5 to -4 <---- from vertex to get a focus of (1,1)

OpenStudy (anonymous):

yup

OpenStudy (firejay5):

then you would subtract -4 - 5 to get -9; which is the directrix

OpenStudy (anonymous):

the dirextix is a line, namely \(x=-9\) but yeah essentially

OpenStudy (firejay5):

okay so thanks for the help; we just got introduced to this today so I was kind of iffy on it

OpenStudy (anonymous):

looks like you got a decent handle on it yw